How many gallons of 70% antifreeze must be mixed with 120 gallons of 20% antifreeze to get a mixture that is 40% antifreeze?
Answer by stanbon(75887)

How many gallons of 70% antifreeze must be mixed with 120 gallons of 20% antifreeze to get a mixture that is 40% antifreeze?
----------------------
Equation:
alcohol + alcohol = alcohol
0.70x + 0.20(120) = 0.40(x+120)
Multiply thru by 100 to get:
70x + 20*120 = 40x + 40*120
30x = 20*120
x = 20*4
x = 80 gallons (amt. of 70% antifreeze needed for the mixture)
=================================
Cheers,
Stan H.